OpenStudy (anonymous):

solve for x ,tan inverse 2x+tan inverse 3x =pi/4

OpenStudy (anonymous):

so you got arctan(2x) + arctan(3x) = pi/4?

OpenStudy (anonymous):

yes.

hero (hero):

\(\tan^{-1}(2x) + \tan^{-1}(3x) = \frac{\pi}{4}\)

OpenStudy (anonymous):

yes. i got it. thanks.

hero (hero):

I can help you with solutions to these, but it would take a while

OpenStudy (anonymous):

no thanks. hero . i have done it. please help me on my next move. i mean next question.